Hi the following question is a 4 part question, I did the first part so I just need confirmation that it was correct and then I need help with parts 2-4 as I am somewhat confused. Thank you in advance!

P.S for 3-C the options are as follows

equals, greater than or less than

will or will not

should or should not

Thanks!

Frieden Companys contribution format income statement for last month is shown below: Sales (45,000 units Variable expenses $1,800,000 1,260,000 Contribution margin Fixed expenses 540,000 432,000 Operating income $108,000 Competition is intense, and Frieden Companys profits vary considerably from one year to the next. Management is exploring opportunities to increase profitability.
Answer #1

Solution

Frieden Company

  1. Contribution margin income statement:

Frieden Company

Contribution Margin Income Statement

Present

Proposed

Amount

Per Unit

Percent

Amount

Per Unit

Percent

Sales

$1,800,000

$40

100%

$1,800,000

$40

100%

Variable Expenses

$1,260,000

$28

70%

$720,000

$16

40%

Contribution Margin

$540,000

$12

30%

$1,080,000

$24

60%

Fixed Expenses

$432,000

$972,000

Operating Income

$108,000

$108,000

Note:

The increase in fixed expenses is by $540,000. Hence, the total fixed expenses for the proposed operations would be $432,000 + $540,000 = $972,000.

  1. Computations of degree of operating leverage, break-even point in dollars, margin of safety in dollars and margin of safety in percentages:

Present

Proposed

a.

Degree of Operating Leverage

5

10

b.

Break-Even Sales in Dollars

$1,440,000

$1,620,000

c.

Margin of Safety in Dollars

$360,000

$180,000

Margin of Safety Percentage

20%

10%

Degree of operating leverage (DOL) –

Degree of operating leverage = Contribution margin/operating income

Present –

Contribution margin = $540,000

Net income = $108,000

DOL = 540,000/108,000 = 5

Proposed –

Contribution margin = $1,080,000

Net income = $108,000

DOL = 540,000/108,000 = 10

Break-even point in dollars (BEP)

Break-even point in dollars = Fixed cost/Contribution margin ratio

Present –

Fixed cost = $432,000

CM ratio = 30%

BEP = 432,000/30% = $1,440,000

Proposed –

Fixed cost = $972,000

CM ratio = 60%

BEP = 972,000/60% = $1,620,000

Margin of Safety in dollars and Percentage –

MOS = Actual Sales – Break-Even Sales in Dollars

Present –

Actual Sales = $1,800,000

BEP Sales = $1,440,000

MOS = 1,800,000 – 1,440,000 = $360,000

MOS percentage = MOS sales/actual sales

Present –

MOS Sales = $360,000

Actual sales = $1,800,000

MOS % = 360,000/1,800,000 = 20%

Proposed –

Actual Sales = $1,800,000

BEP Sales = $1,620,000

MOS = 1,800,000 – 1,620,000 = $180,000

MOS percentage = MOS sales/actual sales

Present –

MOS Sales = $180,000

Actual sales = $1,800,000

MOS % = 180,000/1,800,000 = 10%

3.

3a. Computation of unit sales at which the company would be indifferent:

Unit Sales

45,000

units per month

At The point of indifference,

(Present Contribution margin per unit x unit sales) – Present Fixed Costs = (Proposed Contribution margin per unit x unit sales) – Proposed Fixed Cost

Assuming the unit sales to be A,

($12 x A) - $432,000 = ($24 x A) - $972,000

$24A - $12A = $972,000 - $432,000

$12A = $540,000

A, unit sales = $540,000/$12 = 45,000 units per month

3b. No, should not proceed with the proposed operations

3c. In this case, the indifference point is EQUAL to the current level of sales at which point the upgrade WILL NOT have an impact on the operating income. So, Frieden’s SHOULD NOT proceed to upgrade.

4.

4a. operating income –

Operating Income

$125,000

Increase

Frieden Company

Contribution Margin Income Statement

Proposed

Amount

Per Unit

Percent

Unit Sales

Sales

$1,980,000

$40

100%

49,500 units

Variable Expenses

$1,386,000

$28

70%

Contribution Margin

$594,000

$12

30%

Fixed Expenses

$469,000

Operating Income

$125,000

Computations –

Increase in sales units 10%, sales unit = 45,000 + 10% of 45,000 = 49,500 units per month

No change in Sales price, variable cost, hence CM would be same at $12 per unit

Fixed costs increase by $37,000, = $432,000 + $37,000 = $469,000

Increase in Operating Income = $125,000 - $108,000 = $17,000

4b. Yes

Explanation: The proposed increase in operations with an increase in advertising expense of $37,000 would increase operating income by $17,000. Hence, proceed with the proposed operations.